# Nonlocal mass creation by metricized accelerations in Euclidean matterspace

### v1

@@ Abstract
Keywords Euclidean continuum, adaptive time, field monism, self-accelerating reality, geometrodynamics, non-local self-assembly
Mathematics Subject Classification 00A79, 83C99, 83E05
1 Introduction to field monism
@@
In Europe, the idea of describing nature in terms of the properties of material space-plenum or ponderable ether goes back to the worldview of the ancient Greeks. “Plato in the Timaeus says that matter and space are the same” underlined Aristotle in his Physics [1]. Descartes’s matter-extension [2] from the seventeenth century was supported hundred years later by Lomonosov in his ether-based gravity due to local pushes by material liquid with invisible densities [3]. The pondarable space with inertial heat and ethereal masses was first used quantitatively by Umov in 1873 for his monistic cosmos of continuous energy flows, which can be reinterpreted by secondary relations with volumetric masses in the Newtonian empty space but not vice versa [4].
@@
In Germany, Gustav Mie persistently supported the field theory of matter since his first publication on the continuous elementary particle [5]. It was in 1912, before the emerging era of quantum distributions in the non-local microworld. Einstein accepted the ponderable ether publicly in 1920 [6]. In 1929, the New Yorker Magazine popularized Einstein by noting “People slowly accustomed themselves to the idea that the physical state of space itself were finite physical reality”. For a long time, the monistic wave function for material densities of quantum particles was traditionally associated with the so-called microworld. The mechanical picture of the macroworld relied on Newtonian dualism for localized bodies and continuous forces (or massless fields) in empty space. This dualistic consept dominated at least until non-locality of superconducting currents was established in the laboratory and reliable measurements of megaparsec correlations in astronomy [7, 8].
@@
In 1914, Einstein [9] proposed to nullify the geodesic four-acceleration $a_{\mu}\equiv u^{\nu}\nabla_{\nu}u_{\mu}\equiv u^{\nu}(\nabla_{\nu}u_{\mu}-\nabla_{\mu}u_{\nu})=u^{\nu}(\partial_{\nu}u_{\mu}-\partial_{\mu}u_{\nu})$ for the free fall of the probe body. By convention, the latter does not contribute to the external fields and the local derivatives $\nabla_{\nu}u_{\mu}\equiv\partial_{\nu}u_{\mu}-\Gamma_{\nu\mu}^{\rho}u_{\rho}$ with the symmetric affine correlations $\Gamma_{\nu\mu}^{\rho}=\Gamma_{\mu\nu}^{\rho}$. In the twenties century, the revolution 4-acceleration $a^{\nu}=g^{\mu\nu}a_{\mu}$ did not shake the particle-field dualism of matter. The point is that the relativistic mechanics and metric fields of Einstein, as well as the monistic energy transport of Umov, were satisfactorily referenced by the Newtonian formalism for many testable applications. But why does the dense mass of probe bodies not affect the external field generated universally by all masses? In this paper, we try to reconsider the relativistic four-vector $a^{\mu}$ as a primary notion in the mass-generation mechanism. But how to build geometrodynamics of inertial spaces with locally induced masses by metrically correlated accelerations in order to quantitatively describe nonlocal self-organization of mechanical macrodistributions with negligible dissipation losses?
@@
In summer and autumn of 1915 Hilbert counseltad Einstein regarding the dual action with the classical Lagrangian for inertial sources and the Ricci scalar density for metric fields (around these sources). As a result, the Einstein Equation for ten tensor balances between metric fields and massive densities [10, 11] remained consistent with the Newtonian dualism for massive bodies and their forces or massless fields. In this dualistic approach to the monistic (quantum) reality, negative energies of gravitational or metric fields also remained as the independent notion, which ultimately led to the metric singularity and many mathematical surprises of post-Newtonian physics.
@@
Nowadays the Cartesian mechanics of inertial space seems to have its renaissance due to new experimental data, including astro-correlations [7, 8]. This paper reveals covariant updates for the static equilibrium of the extended mass-energy, where the Ricci scalar density $R(r)\equiv(\mu_{a}+\mu_{p})8\pi G/c^{2}=4r_{o}^{2}/[r^{2}(r+r_{o})^{2}]$ is proportional to the equal densities of the active and passive (or reactive) masses [12]. The monistic macrophysics of inertial fields or field masses should clarify how inhomogeneous densities of kinetic energy resist gravitational collapse or resolve the Bentley paradox of Newtonian physics. The goal of our flatspace relativism with only four Hilbert variations under the six geometric bounds is to maintain Cartesian worldview quantitatively and to initiate the adaptive geometrodynamics of correlated inertial densities for further applications in nonlocal macrophysics, life sciences, engineering, and novel energo-informatics of big data.
@@ 2 Accelerations define field masses and monistic action
The static self-organization of the nonlocal charge in the laboratory takes the equilibrium energy density $\varphi_{\!{}_{G}}{\sqrt{G}}\mu(r^{\prime}=r)\equiv c^{2}\mu(r^{\prime}=r)$ around the point vertex at the coordinate origin. This static distribution follows the Poisson equation $\mu(r)=-\delta^{ij}\partial_{i}w_{j}/4\pi G$ with the metric field intensity $w_{i}(r)=\delta_{ij}w^{j}=-Gm(x^{j}/r)/[r(r+r_{o})]\equiv-c^{2}{\partial_{i}}ln{\sqrt{g_{oo}(r)}}$, $g_{oo}(r)=r^{2}/(r+r_{o})^{2}=1/g^{oo}(r)$ [12]. But inhomogeneous profiles of extended masses and their rest energies should be stabilized in the classical mechanics by some kind of non-Newtonian forces associated with the so-called Poincaré pressure [13]. How to balance the self-organizing forces in the monistic geometrodynamics of material space with affine connections or nonlocal correlations?
Again, the essential message of the radial solutions in [12] for static inertial distributions was the equivalence of the active (attractive) and passive (reactive) mass-energy densities, $\mu_{a}(r)c^{2}\equiv-c^{2}\partial_{i}w_{j}(r)\delta^{ij}/4\pi G$ $=w_{i}(r)w_{j}(r)\delta^{ij}/4\pi G\equiv\mu_{p}(r)c^{2}$. Both densities contribute to the Ricci scalar $R(r)=(\mu_{a}c^{2}+\mu_{p}c^{2})8\pi/\varphi_{\!{}_{G}}^{2}$ $=4r_{o}^{2}/[r(r+r_{o})]^{2}=2u_{o}i^{o}(r)-2f_{oi}(r)f^{oi}(r)$ in the co-moving frame with $u_{o}={\sqrt{g_{oo}(r)}}\equiv g_{o}(r)=r/(r+r_{o}),g_{i}\equiv g_{oi}/g_{o}=0,g_{oi}=0,g^{oi}=0,g^{ij}=-\delta^{ij}$, $g_{ij}=-\delta_{ij}$, $u_{i}=u^{i}=0$, $f_{oi}=-\partial_{i}g_{o}=-(x_{i}r_{o}/r)/(r+r_{o})^{2},f^{oi}=g^{oo}g^{ij}f_{oj}=\delta^{ij}x_{i}r_{o}/r^{3},(-f_{oi}f^{oi})=r^{2}_{o}/[r(r+r_{o})]^{2}=g_{o}i^{o},i^{o}=r_{o}^{2}/[r^{3}(r+r_{o})]$. The invariant sum of equal static contributions clarifies the tensor structure of the Lagrangian ${\cal L}(x)=-\varphi_{\!{}_{{}_{G}}}^{2}R(x)/{16\pi}$ or the scalar curvature $R(x)=2u_{\nu}(x)i^{\nu}(x)-f_{\mu\nu}(x)f^{\mu\nu}(x)>0$ in the monistic action $S_{m}$ for inertial (massive) fields by the local divergence of the coordinate self-accelerations $a^{\mu}(x)\equiv c^{2}u_{\nu}(x)f^{\nu\mu}(x)$:
@@
By defining the symmetric energy-momentum tensor $T_{\mu\nu}(x)$ from ${\sqrt{-g}}T_{\mu\nu}/2\equiv\partial({\sqrt{-g}}{\cal L})/\partial g^{\mu\nu}$ according to Hilbert [11] and using $\partial{\sqrt{-g}}\equiv-{\sqrt{-g}}g_{\mu\nu}\partial g^{\mu\nu}/2$ one gets the following densities of massive (inertial, material) fields:
@@
$\cases{\frac{{\sqrt{-g}}}{2}T_{\mu\nu}\equiv\frac{c^{4}{\sqrt{-g}}}{16\pi G}(g_{\mu\mu}\nabla_{\lambda}a^{\lambda}-\nabla_{\mu}a_{\nu}-\nabla_{\nu}a_{\mu}),\cr T^{\nu}_{\mu}\equiv g^{\nu\lambda}T_{\mu\lambda}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(\delta_{\mu}^{\nu}\frac{{B}}{2}-{B}_{\mu}^{\nu}),\ B_{\mu\nu}\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu},\cr{B}^{\nu}_{\mu}\equiv g^{\nu\lambda}B_{\mu\lambda},B^{\mu}_{\mu}\equiv 2\nabla_{\mu}a^{\mu}=R_{\mu}^{\mu}\Rightarrow\frac{4r_{o}^{2}}{r^{2}(r+r_{o})^{2}},a^{o}=0,u^{i}=0,\cr T^{\mu}_{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{\mu}a^{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{o}a^{o}\!+\!\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{i}a^{i}\equiv(\mu_{p}\!+\!\mu_{a})c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(B_{o}^{o}\!+\!B^{i}_{i}),\cr\mu_{p}c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{o}a^{o}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}(\partial_{o}a^{o}+\Gamma_{o\lambda}^{o}a^{\lambda})\Rightarrow\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\Gamma_{oi}^{o}a^{i}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}r_{o}^{2}}{4\pi r^{2}(r+r_{o})^{2}},\cr\mu_{a}c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{i}a^{i}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}(\partial_{i}a^{i}+\Gamma_{i\lambda}^{i}a^{\lambda})\Rightarrow\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\partial_{i}a^{i}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}r_{o}^{2}}{4\pi r^{2}(r+r_{o})^{2}}.}$ (4)

### v2

@@ Abstract
Keywords Classical Theories of Gravity, Space-Time Symmetries, Gauge Symmetry
1 Introduction
1.1 Monism versus Dualism
@@
In Europe, the concept of describing nature in terms of the properties of a material space-plenum or ponderable ether dates back to the worldview of the ancient Greeks. “Plato in the Timaeus says that matter and space are the same” underlined Aristotle in his Physics [1]. Descartes’s matter-extension [2, 3], from the seventeenth century, was supported a hundred years later by Lomonosov with his ether-based gravity theory, which suggested local pushes by a material liquid with invisible densities [4]. The pondarable space with inertial heat and etheric masses was first described quantitatively by Umov’s continuous transport of energy flows in a monistic cosmos-medium [5]. This monistic distribution of continuous mass-energy densities can be reinterpreted and approximated by secondary (as Umov underlined) relations with volumetric masses in the Newtonian empty space, but not vice versa.
@@
In Germany, Gustav Mie persistently supported the field theory of matter since his first publication on the continuous elementary particle [6] in 1912, before the emerging era of quantum distributions in the non-local microworld. Einstein publicly accepted the ponderable ether in 1920 [7]. In 1929, the New Yorker Magazine popularized Einstein by noting “People slowly accustomed themselves to the idea that the physical states of space itself were the finite physical reality”. For a long time, the monistic wave function for material densities of quantum particles was traditionally associated with the so-called microworld. The mechanical picture of the macroworld relied on Newtonian dualism for localized bodies and continuous forces (or massless fields) in empty space. This dualistic consept dominated until non-locality of superconducting currents was established in the low-temperature laboratory. In this century, astronomers are publishing more and more about megaparsec correlations [8, 9].
1.2 Did Einstein geometrize the void or the massive field?
@@
In 1914, Einstein [10] proposed to nullify the geodesic four-acceleration $a_{\mu}\equiv u^{\nu}\nabla_{\nu}u_{\mu}\equiv u^{\nu}(\nabla_{\nu}u_{\mu}-\nabla_{\mu}u_{\nu})=u^{\nu}(\partial_{\nu}u_{\mu}-\partial_{\mu}u_{\nu})$ for the free fall of the probe body. By convention, the latter does not contribute to the external fields and the local derivatives $\nabla_{\nu}u_{\mu}\equiv\partial_{\nu}u_{\mu}-\Gamma_{\nu\mu}^{\rho}u_{\rho}$ with the symmetric affine connections (correlations) $\Gamma_{\nu\mu}^{\rho}=\Gamma_{\mu\nu}^{\rho}$. In the twenties century, the revolution 4-acceleration $a^{\nu}=g^{\mu\nu}a_{\mu}$ did not shake the particle-field dualism of matter. The point is that the relativistic mechanics and metric fields of Einstein, as well as the monistic energy transport of Umov, were satisfactorily referenced by the Newtonian formalism for many testable applications.
@@
In summer and autumn of 1915 Hilbert [11] counseltad Einstein regarding the dual action with the classical Lagrangian for inertial sources and the Ricci scalar density for metric fields (around these sources). As a result, the Einstein Equation for ten tensor balances between metric fields and massive densities [12] remained consistent with the Newtonian dualism for massive bodies and their forces or massless fields. In this dualistic approach to the monistic (quantum) reality, negative energies of gravitational or metric fields also remained as the independent notion, which ultimately led to the metric singularity and many mathematical surprises of post-Newtonian physics.
For physical theories, practice is indeed crucial. But observations and measurements can only falsify a naive concept, not select the right one among competing theoretical constructions [13]. The dual partitioning of matter into massive particles and massless force fields in the Newtonian worldview leads to many computational absurdities, including the divergence of energy for a singularity. In 1939 Einstein completed the thought experiment [14] in order to reject Schwarzschild’s metric singularities for realistic kinds of mechanical movements - “… at least not faster than light”.
Why does the dense mass of probe bodies not affect the external field generated universally by all masses? Many dualistic shortcomings have encouraged researchers to reconsider sparse and dense energy regions in monistic terms of continuous mass. For example, “A thrown stone is, from this point of view, changing field, where the states of greatest field intensity travel through space with the velocity of the stone” Einstein and Infeld whrote about the non-dual (monistic) reinterpretation of the observable reality [15]. We tend to maintain the relativistic four-vector $a^{\mu}$ as a primary notion in the mass-generation mechanism. But how to build auto-dynamics of inertial spaces with locally induced masses by metrically correlated accelerations in order to quantitatively describe nonlocal self-organization of mechanical macrodistributions with negligible dissipation losses?
1.3 Macroscopic entanglement revives material space-plenum
@@
Nowadays the Cartesian mechanics of inertial space is experiencing a renaissance due to new experimental data, including astro-correlations [8, 9]. This paper presents covariant updates for the static equilibrium of extended mass-energy, where the Ricci scalar density $R(r)\equiv(\mu_{a}+\mu_{p})8\pi G/c^{2}$ $=4r_{o}^{2}/[r^{2}(r+r_{o})^{2}]$ is proportional to the equal densities of active and passive (or reactive) masses [16]. The monistic macrophysics of inertial fields or field masses aims to clarify how inhomogeneous densities of kinetic energy resist gravitational collapse or resolve the Bentley paradox of Newtonian physics. The goal of our flatspace relativism, with only four Hilbert variations under the six geometric constraints or inherent symmetries of space-time geometry, is to maintain Cartesian worldview with its matter-extension quantitatively. Additionally, we aim to initiate the adaptive auto-dynamics of metrically correlated inertial densities (called below geometrodynamics) for further applications in nonlocal macrophysics, life sciences, engineering, and the novel energo-informatics of big data.
@@ 2 Accelerations define field masses and monistic action
2.1 Pseudo-Riemann definitions of material space with its local velocities and inertial currents
@@
The static self-organization of the nonlocal charge in the laboratory achieves equilibrium energy density $\varphi_{\!{}_{G}}{\sqrt{G}}\mu(r^{\prime}=r)\equiv c^{2}\mu(r^{\prime}=r)$ around the point vertex at the coordinate origin. This static distribution follows the Poisson equation $\mu(r)=-\delta^{ij}\partial_{i}w_{j}/4\pi G$ with the metric field intensity $w_{i}(r)=\delta_{ij}w^{j}=-Gm(x^{j}/r)/[r(r+r_{o})]\equiv-c^{2}{\partial_{i}}ln{\sqrt{g_{oo}(r)}}$, $g_{oo}(r)=r^{2}/(r+r_{o})^{2}=1/g^{oo}(r)$ [16]. However, inhomogeneous profiles of extended masses and their rest energies require stabilization by non-Newtonian forces, such as those related to the so-called Poincaré pressure [17]. How can we balance the self-organizing forces in the monistic geometrodynamics of material space with affine connections (used in mathematics) or nonlocal correlations (used in physics)?
2.2 Monistic action for nonlocal masses with metrically correlated four-currents
The essential message of the radial solutions in [16] for static inertial distributions was the equivalence of the active (attractive) and passive (reactive) mass-energy densities, $\mu_{a}(r)c^{2}\equiv-c^{2}\partial_{i}w_{j}(r)\delta^{ij}/4\pi G$ $=w_{i}(r)w_{j}(r)\delta^{ij}/4\pi G\equiv\mu_{p}(r)c^{2}$. Both densities contribute to the Ricci scalar $R(r)=(\mu_{a}c^{2}+\mu_{p}c^{2})8\pi/\varphi_{\!{}_{G}}^{2}$ $=4r_{o}^{2}/[r(r+r_{o})]^{2}=2u_{o}i^{o}(r)-2f_{oi}(r)f^{oi}(r)$ in the co-moving frame with $u_{o}={\sqrt{g_{oo}(r)}}\equiv g_{o}(r)=r/(r+r_{o}),g_{i}\equiv g_{oi}/g_{o}=0,g_{oi}=0,g^{oi}=0,g^{ij}=-\delta^{ij}$, $g_{ij}=-\delta_{ij}$, $u_{i}=u^{i}=0$, $f_{oi}=-\partial_{i}g_{o}=-(x_{i}r_{o}/r)/(r+r_{o})^{2},f^{oi}=g^{oo}g^{ij}f_{oj}=\delta^{ij}x_{i}r_{o}/r^{3},(-f_{oi}f^{oi})=r^{2}_{o}/[r(r+r_{o})]^{2}=g_{o}i^{o},i^{o}=r_{o}^{2}/[r^{3}(r+r_{o})]$. The invariant sum of equal static contributions clarifies the tensor structure of the Lagrangian ${\cal L}(x)=-\varphi_{\!{}_{{}_{G}}}^{2}B(x)/{16\pi}$ with the scalar density $B(x)=2u_{\nu}(x)i^{\nu}(x)-f_{\mu\nu}(x)f^{\mu\nu}(x)>0$ in the monistic action $S_{m}$ for inertial (massive) fields by the local divergence of the coordinate self-accelerations $a^{\mu}(x)\equiv u_{\nu}(x)f^{\nu\mu}(x)$:
@@ 2.2 Monistic action for nonlocal masses with metrically correlated four-currents
By defining the monistic field tensor $T_{\mu\nu}(x)$ from the Hilbert formula for the substance stress-energy tensor [11] ${\sqrt{-g}}T_{\mu\nu}/2\equiv\partial({\sqrt{-g}}{\cal L})/\partial g^{\mu\nu}$ with the opposite sigh (to match the dualistic balance of Hilbert fields and gravitating substances) and using $\partial{\sqrt{-g}}\equiv-{\sqrt{-g}}g_{\mu\nu}\partial g^{\mu\nu}/2$ for the metric space-time, one gets the following notations for inertial densities of continuous matter in monistic field theory:
@@
$\cases{\frac{{\sqrt{-g}}}{2}T_{\mu\nu}\equiv\frac{c^{4}{\sqrt{-g}}}{16\pi G}(g_{\mu\mu}\nabla_{\lambda}a^{\lambda}-\nabla_{\mu}a_{\nu}-\nabla_{\nu}a_{\mu}),\cr T^{\nu}_{\mu}\equiv g^{\nu\lambda}T_{\mu\lambda}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(\delta_{\mu}^{\nu}\frac{{B}}{2}-{B}_{\mu}^{\nu}),\ B_{\mu\nu}\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu},\cr{B}^{\nu}_{\mu}\equiv g^{\nu\lambda}B_{\mu\lambda},B^{\mu}_{\mu}\equiv 2\nabla_{\mu}a^{\mu}\Rightarrow\frac{4r_{o}^{2}}{r^{2}(r+r_{o})^{2}},a^{o}=0,u^{i}=0,\cr T^{\mu}_{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{\mu}a^{\mu}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{o}a^{o}\!+\!\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{i}a^{i}\equiv(\mu_{p}\!+\!\mu_{a})c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi}(B_{o}^{o}\!+\!B^{i}_{i}),\cr\mu_{p}c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{o}a^{o}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}(\partial_{o}a^{o}+\Gamma_{o\lambda}^{o}a^{\lambda})\Rightarrow\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\Gamma_{oi}^{o}a^{i}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}r_{o}^{2}}{4\pi r^{2}(r+r_{o})^{2}},\cr\mu_{a}c^{2}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\nabla_{i}a^{i}\equiv\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}(\partial_{i}a^{i}+\Gamma_{i\lambda}^{i}a^{\lambda})\Rightarrow\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\partial_{i}a^{i}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}r_{o}^{2}}{4\pi r^{2}(r+r_{o})^{2}}.}$ (4)
The Einstein-type tensor $g_{\mu\nu}(B^{\lambda}_{\lambda}/2)-B_{\mu\nu}\equiv 8\pi T_{\mu\nu}/{\varphi_{\!{}_{{}_{G}}}^{2}}$ and its tensor curvature $B_{\mu\nu}\equiv\nabla_{\mu}a_{\nu}+\nabla_{\nu}a_{\mu}$ $=R_{\mu\nu}+\theta_{\mu\nu}$, where $R_{\mu\nu}\equiv\partial_{\lambda}\Gamma^{\lambda}_{\mu\nu}-\partial_{\nu}\Gamma^{\lambda}_{\mu\lambda}+\Gamma^{\lambda}_{\mu\nu}\Gamma^{\rho}_{\rho\lambda}-\Gamma^{\lambda}_{\mu\rho}\Gamma^{\rho}_{\nu\lambda}$, $g^{\mu\nu}\theta_{\mu\nu}\equiv 0$, $g^{\mu\nu}B_{\mu\nu}=g^{\mu\nu}R_{\mu\nu}\equiv R$ and $B_{\mu\nu}\neq R_{\mu\nu}$, are based in our novel analyses on the covariant derivatives of four-accelerations $a_{\mu}$. Hereinafter we require strict 3D flatness, $\kappa_{ij}\equiv g_{i}g_{j}-g_{ij}\equiv\delta_{ij}$, for the monistic continuum of field masses. The corresponding six constraints for Euclidean 3-space, usually required by theorists for electrodynamics and quantum mechanics, have not yet been falsified in the available experiments. The metric component $g_{oo}=r^{2}/(r+r_{o})^{2}$ for the radial self-organisation of static material continuum [16] is free from pecularities. Such metric field allows to recalculate the main predictions of the dualistic General Relativity (gravitational light bending, Shapiro ekho delay, Merciry perehelion precession) in terms of flat space with adaptively delayed local time.
The stress-energy tensor $T_{\mu\nu}(x)\equiv g_{\nu\lambda}(x)T^{\lambda}_{\mu}(x)$ elucides the local generation of non-local (field) masses, highlighting the non-Newtonian monism of material reality and invariant fields pivotal in modern metric theories. Their geometric interplay bridges classical and quantum physics, inspiring cosmological models and advancing geometrodynamic paradigms.
2.3 Inherent space-time symmetris for Euclidean 3-geometry
The Euclidean constraints $\kappa_{ij}=\delta_{ij},g^{ij}=\delta^{ij}$ can be called the inherent symmetries of the space-time geometry in the basic geometrodynamic equality $u_{\mu}u^{\mu}\equiv 1$. These symmetries are maintained by the adaptive local time $d\tau\equiv g_{\mu}dx^{\mu}/c\neq dx^{o}/c$, where $g_{\mu}\equiv g_{o\mu}/g_{o}$. Besides beying a function of metric and motion changes, the local time for physical processes is not a primary concept in the correlated field theory for a closed metric (hierachial) system. It is a secondary variable that depends on energy-momentum variations, but not vice versa. However, the common shift $dx^{o}/c$ of the universal (world, Newtonian) parameter $x^{o}/c$ can be considered as a primary notion, like three Euclidian (universal) coordinates $x^{i}$ for the Cartesian matter-extension. Essentially, this implies that four temporal components $g_{o\mu}$ together with the metric 4-potential $g_{\mu}c^{2}$ and the corresponding time rate (experienced locally by a probe body) are consistent with observable consequences. Time and metric variations are not a primary cause of observable phenomena, but a nonlocal consequence of the correlated kinetics of all massive fields.